((x-2)/(x-1))+((x-3)/(2x-5))=1
I was told to solve this equation but im having a heck of a time. The whole =1 thing is killin me and throwing me way off. Please help if you can.

Answer by venugopalramana(3286)

((x-2)/(x-1))+((x-3)/(2x-5))=1
LCM IS (X-1)(2X-5)..SO MULTIPLY THROUGHOUT WITH THE LCM ..WE GET..
(x-2)(2x-5)+(x-3)(x-1)=(x-1)(2x-5)
2x^2-5x-4x+10+x^2-x-3x+3=2x^2-5x-2x+5
2x^2-5x-4x+10+x^2-x-3x+3-2x^2+5x+2x-5=0
x^2-6x+8=0
x^2-4x-2x+8=0
x(x-4)-2(x-4)=0
(x-2)(x-4)=0
x-2=0....or....x-4=0
x=2....or.....4
you can check back by substitution